[F] Eun-Lee is flying from Calgary to Vancouver when she has to make an emergency landing in the mountains near Lake Louise. She has can either land on the west side of the lake, where there is a 1100m rock path (coefficient of friction between the tires nd rock=0.22). . Or she can land on the east side of the lake, where there is a 850 m grass path (coefficient of friction between the tires and grass is 0.44). She will be going 250 km/h * r when she hits the ground. She knows her brakes are not going to work. Which side of the lake should she land on to save the day?
1
Expert's answer
2019-11-18T10:36:41-0500

1)


"a_1=\mu _1g"

"V^2-v^2=2a_1s_1"

"\left(\frac{250}{3.6}\right)^2-v^2=2(0.22)(9.8)(1100)"

The speed after this path is

"v=8.9\frac{m}{s}"

2) In this case


"V^2<2a_2s_2"

"\left(\frac{250}{3.6}\right)^2<2(0.44)(9.8)(850)\to 4823<7330"

Thus, she stops on trhis path.


Answer: the west side of the lake.